OpenStudy (anonymous):

cos(x+y) =4/5,sin(x-y)=5/13 (x+y), (x-y) are acute..... Find tan2x?

OpenStudy (anonymous):

Do you have to find \(\tan2x\) or \(\tan^2x\)?

OpenStudy (anonymous):

\[\begin{align*}\tan2x&=\tan (x+x)\\ &=\tan\bigg((x+y)+(x-y)\bigg)\\ &=\frac{\tan(x+y)+\tan(x-y)}{1-\tan(x+y)\tan(x-y)}\\\\ &=\frac{\dfrac{\sin(x+y)}{\cos(x+y)}+\dfrac{\sin(x-y)}{\cos(x-y)}}{1-\dfrac{\sin(x+y)}{\cos(x+y)}\cdot\dfrac{\sin(x-y)}{\cos(x-y)}}\\ \end{align*}\]

OpenStudy (anonymous):

Given that \(x+y\) and \(x-y\) are acute, both angles are in the first quadrant, which mean their sines and cosines are positive. If \(\cos(x+y)=\dfrac{4}{5}\), then you can easily find \(\sin(x+y)\): \[\begin{align*}\cos^2(x+y)+\sin^2(x+y)&=1\\ \dfrac{16}{25}+\sin^2(x+y)&=1\\ \sin(x+y)&=\sqrt{1-\frac{16}{25}} \end{align*}\] (you take the positive root because you know sine must be positive) You can find \(\cos(x-y)\) in a similar fashion.

OpenStudy (anonymous):

How elegant the solution is!!